At present Allen is 10 years younger than Scott. Ten years from now, Scott’s age will be 3 times the present age of Allen. Find the present age of Allen and Scott.
Found 2 solutions by addingup, ikleyn:
Answer by addingup(3677)

A = S-10
A+10 = 3A
S-10+10 = 3(S-10)
S = 3S-30
-2S = -30
S = 15
and A = 15 - 10 = 5
~~~~~~~~~~~~~~~~~~~~~
Check:
A+10 = 3A
5+10 = 3(5)
15 = 15 Correct

Answer by ikleyn(52776)

.

            Since the solution by @addingup is  ABSOLUTELY WRONG,  I came to bring the correct solution.


A = S-10      (1)
S+10 = 3A     (2)


Substitute expression  A = S-10  from eq(1) into eq(2), replacing A in equation (2).


S + 10 = 3*(S-10)

S + 10 = 3S - 30

10 + 30 = 3S - S

2S = 40   ====>  S = 40/2 = 20.


ANSWER.  Scott is 20 years now.  Allen is 20-10 = 10 years now.


Check.   Eq(1):  A = S-10:   10 = 20-10      ! Correct !

         Eq(2):  S+10 = 3A:  20+10 = 3*10    ! Correct !

Solved and checked.